OpenStudy (anonymous):

Pies are ordered from the Pizza Palace according to their diameter. Large:sixteen inches Medium:fourteen inches Small:twelve inches How many more square inches are in the large pizza than in the small pizza? @ajanijones

OpenStudy (ajanijones):

Nova the last question is 16 not 32.

OpenStudy (ajanijones):

16-12= 4

OpenStudy (ajanijones):

So 4in

OpenStudy (anonymous):

thnx dude ur awesome

OpenStudy (anonymous):

Square inches indicates Area \[A = \pi r ^{2}\] halve the diamaters of the large pizza and the small pizza and find their area

OpenStudy (ajanijones):

Because you just subtract so 16-12=?

OpenStudy (anonymous):

no ajani, you need the areas

OpenStudy (ajanijones):

No its not area read the question again.

OpenStudy (anonymous):

their diameter is 16in and 12 in. they're asking the difference between their square inches.

OpenStudy (anonymous):

the answer that you @ajanijones wasnt there

OpenStudy (ajanijones):

N0

OpenStudy (anonymous):

@CliffSedge Please read this question. it's an Area question correct?

OpenStudy (anonymous):

sorry the answers are twelve point five six square inches forty-seven point one zero square inches eighty-seven point nine two square inches three hundred fifty-one point six eight square inches

OpenStudy (anonymous):

\[A = \pi 8^{2} - A = \pi 6^{2}\]

OpenStudy (anonymous):

@ajanijones u there dude

OpenStudy (anonymous):

200.96 - 113.04 = 87.92 square inchess. Answer is C.

OpenStudy (anonymous):

so what do u think it is @ECSarabia

OpenStudy (anonymous):

C. After doing the math.

OpenStudy (anonymous):

thnx